I need help on an algebra 1 question by solving systems using substitution. Here is the question:
Michelle and Pam are running a 200-m race. Michelle runs at an average of 7.5 m/s. Pam averages 7.8m/s, but she starts 1 s after Michelle.
A. how long will it take Pam to catch up to Michelle?
B. Will Pam overtake Michelle before the finish line?
I have already come up with an equation.
7.5x=200
7.8(x-1)=200
I attempted to answer question A, though I'm pretty sure i hadn't done it right.
7.5=7.8-7.8
-0.3x=-7.8
x=7.5

Answer by ankor@dixie-net.com(22740)

Michelle and Pam are running a 200-m race.
Michelle runs at an average of 7.5 m/s. Pam averages 7.8m/s, but she starts 1 s after Michelle.
:
A. how long will it take Pam to catch up to Michelle?
Let t = time required for P to catch M.
then
(t+1) = travel time for M when she is caught
When this happens, they will have traveled the same distance
Write a distance equation; dist = speed * time
7.8t = 7.5(t+1)
7.8t = 7.5t + 7.5
7.8t - 7.5t = 7.5
.3t = 7.5
t = 7.5/.3
t = 25 seconds for P to catch M
:
B. Will Pam overtake Michelle before the finish line?
Find the distance they have covered when P catches M
25 * 7.8 = 195 m
26 * 7.5 = 195 m
Yes, 5 m before the end of the race